OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (3, 4) and (5, 16)?

TheSmartOne (thesmartone):

\(\bf\huge~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @busta34 :)

TheSmartOne (thesmartone):

Do you know what the slope-intercept form is?

TheSmartOne (thesmartone):

And do you know how to calculate the slope? :)

OpenStudy (amistre64):

solve for y:\[\frac{y-4}{x-3}=\frac{y-16}{x-5}\]

TheSmartOne (thesmartone):

Hmm, couldn't we just calculate the slope since we are given the two points. And then plug in the slope into the slope-intercept format and then plug in a point (any of the two points) and calculate the y-intercept?

OpenStudy (amistre64):

we could if we wanted a boring way :)

OpenStudy (amistre64):

why bother doing parts and trying to keep track of them ... thats just barbaric ... the new math is sleek, impressive, and itll wow your friends and teachers.

TheSmartOne (thesmartone):

Haha :)

OpenStudy (anonymous):

I don't really understand what the question was asking.. about slope intercept form

TheSmartOne (thesmartone):

Slope intercept form is \(\sf y=mx+b\) Where m is the slope, and b is the y-intercept.

OpenStudy (amistre64):

\[\frac{y-4}{x-3}=\frac{y-16}{x-5}\] \[\frac{x-5}{x-3}=\frac{y-16}{y-4}\] \[\frac{x-3-2}{x-3}=\frac{y-4-12}{y-4}\] \[1-\frac{2}{x-3}=1-\frac{12}{y-4}\] \[\frac{1}{x-3}=\frac{6}{y-4}\] \[y-4=6(x-3)\] etc ...

OpenStudy (amistre64):

oy the lag .....

OpenStudy (anonymous):

much prefer slope intersect method